# Toeplitz Operators Defined between Köthe Spaces

### v1

@@ 5. Toeplitz Operators Defined between Power Series Spaces
Denoting the operator defined by the Toeplitz matrix as $T_{\theta}$, in this case one can write this operator as a sum of two operators defined by a lower and an upper Toeplitz matrix, that is,
@@
Let $\alpha$ and $\beta$ be a positive sequences that tend to infinity. Assume that $\alpha$ satisfies the condition in 3.4. $T_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is continuous if and only if $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ce^{-\frac{1}{m}\alpha_{n}}\quad\quad\forall n\geq k.$ (5.1)
@@
Let $\alpha$ and $\beta$ be a positive sequences that tend to infinity. Assume that $\alpha$ satisfies the condition in 3.4 and $\Lambda_{\infty}(\beta)$ is nuclear. $T_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is compact if and only if $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n}}\quad\quad\forall n\geq k.$ (5.2)
Proof.
This follows from Theorem 3.3 and Theorem 4.6. ∎
Remark 5.3.
Remark 5.4.
The family of the Toeplitz operators $T_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ for $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ is $\tilde{I}$-tame from Proposition 3.4 and Proposition 4.7 where $\tilde{I}(n)=Mn$ for every $n\in\mathbb{N}$ and $M$ in the condition 3.4.
Theorem 5.5.
Let $\beta$ be a positive sequence that tends to infinity and satisfies the condition in 3.4. $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\beta)$ is continuous if and only if $\theta\in\Lambda_{\infty}(\beta)\oplus(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ce^{m\alpha_{n}}\quad\quad\forall n\geq k.$ (5.3)
Proof.
This follows from Theorem 3.5 and Theorem 4.2. ∎
Theorem 5.6.
Let $\beta$ be a positive sequence that tends to infinity and satisfies the condition in 3.4. $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\beta)$ is compact if and only if $\theta\in\Lambda_{\infty}(\beta)\oplus(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ce^{m\alpha_{n}}\quad\quad\forall n\geq k.$ (5.4)
@@
Remark 5.8.
The family of the Toeplitz operators $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ for $\theta\in\Lambda_{\infty}(\beta)\oplus(\Lambda_{\infty}(\beta))^{\prime}$ satisfying inequality 3.4 is $\tilde{I}$-tame from Proposition 3.8 and Proposition 4.4 where $\tilde{I}(n)=\max(M,2)n$ for every $n\in\mathbb{N}$ and M in the condition 3.4 .
@@
Let $\Lambda_{\infty}(\alpha)$ be any power series space of infinite type and $\Lambda_{1}(\beta)$ be a nuclear power series spaces of finite type. $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{1}(\beta)$ is continuous and compact if and only if $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{\infty}(\beta))^{\prime}$.
@@
By choosing the identity map $I$ as $S$, one can show that the family of the Toeplitz operators $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{1}(\beta)$ for $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{\infty}(\alpha))^{\prime}$ is $I$-tame.
Remark 5.11.
@@
Acknowledgment
The results in this paper were obtained while the author visited at University of Toledo. The author would like to thank TUBITAK for their support.

### v3

@@ 5. Toeplitz Operators Defined between Power Series Spaces
Therefore every Toeplitz matrix can be associated with two sequences. Denoting the operator defined by the Toeplitz matrix as $T_{\theta}$, one can write this operator as a sum of two operators defined by a lower and an upper Toeplitz matrix, that is,
@@
Let $\alpha$ be a stable sequence. A Toeplitz operator $T_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is well-defined and continuous if $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ce^{-\frac{1}{m}\alpha_{n}}\hskip 43.36243pt\forall n\in\mathbb{N}.$ (5.1)
Proof.
Let assume that Toeplitz matrix associated with $T_{\theta}$ is given by sequnces $\widehat{\theta}$ and $\widecheck{\theta}$. We can choose $\theta^{\prime}_{0}$ and $\theta^{\prime\prime}_{0}$ to be non-zero. Then Toeplitz operator $T_{\theta}$ is defined as
From Theorem (3.3), $\widehat{T}_{\widehat{\theta}}$ is well-defined and continous if and only if $\widehat{\theta}\in\Lambda_{1}(\beta)$ and the condition (5.1) is satisfied. From Theorem 4.6, $\widecheck{T}_{\widecheck{\theta}}$ is well-defined and continous if and only if $\widecheck{\theta}\in(\Lambda_{1}(\alpha))^{\prime}$ and the condition (5.1) is satisfied. Therefore, these give us that $T_{\theta}$ is well-defined and continuous if $\theta=\widehat{\theta}\oplus\widecheck{\theta}\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ and the condition (5.1) is satisfied. ∎
Theorem 5.2.
Let $\alpha$ be a stable sequence. A Toeplitz operator $T_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is compact if a $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ce^{-\frac{1}{m}\alpha_{n}}\hskip 28.90755pt\forall n\geq k.$ (5.2)
Proof.
The proof is similar to the proof of Theorem 5.1 and follows from Theorem 3.4 and Theorem 4.7. ∎
Remark 5.1.
Theorem 5.3.
Let $\beta$ be a stable sequence and $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type. A Toeplitz operator $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\beta)$ is well-defined and continuous if $\theta\in\Lambda_{\infty}(\beta)\oplus(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ce^{m\alpha_{n}}\hskip 21.68121pt\forall n\geq k.$ (5.3)
@@
Let $\beta$ be a stable sequence and $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type. A Toeplitz operator $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\beta)$ is compact if $\theta\in\Lambda_{\infty}(\beta)\oplus(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ce^{m\alpha_{n}}\hskip 21.68121pt\forall n\geq k.$ (5.4)
@@
Let $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type and $\Lambda_{1}(\beta)$ be a nuclear power series spaces of finite type. $T_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{1}(\beta)$ is well-defined, continuous and compact if $\theta\in\Lambda_{1}(\beta)\oplus(\Lambda_{\infty}(\beta))^{\prime}$.
@@
Acknowledgments
The results in this paper were obtained while the author visited at University of Toledo. I would like to thank TUBITAK for their support and Prof. Dr. Sönmez Şahutoğlu for sharing Jasiczak’s paper with me, which enabled me to create this work.